OpenStudy (anonymous):

Graph the parabola: x - 2 = 1/8 (y+1)^2 my attempt to solve: h = 2 p = 2 k = 1 vertex: (2,1) directrix: 0 focus: (4,1) My question is actually - do I take h as 2 or -2? ._. I'm never sure when to ignore the sign and when not to do that. Also, should k be -1 as well?

OpenStudy (dumbcow):

k=-1 its x = a(y-k)^2 + h so y+1 = y-(-1)

OpenStudy (anonymous):

Ok thanks. What if you have this one though: y - 2 = 1/16 (x - 1)^2 Is h = -1 or 1?

OpenStudy (dumbcow):

h=1 a trick is to set what is in parenthesis equal to 0 x-1 = 0 x = 1 thus h=1

OpenStudy (anonymous):

I finally got it. Thanks a lot!

OpenStudy (dumbcow):

yw